Length of an arc



Find the length of an arc of a circle having the radius of 8 cm and subtended by a central angle of 330°.


Find the radius of a circle subtended by a central angle of π/4 radius and having an arc length of π.


On a circle of radius 100 cm, the arc intercepted a central angle of 3π/4. What is the arc length?


1
Expert's answer
2021-01-25T04:10:27-0500

a. length of an arc = 2"\pi r(\frac{{1}}{360}\theta)"


"l=" "2 \times \pi \times 8 \times (\frac{{330}}{360}) = 46. 0767"


"l" = 46.0767 cm



b.

"\theta" = "\frac{1}{4}\pi" = 45"^0"

"l=" "2 \times \pi \times r \times (\frac{{45}}{360})"


but "l=\pi"

then

"\pi=" "2 \times \pi \times r \times (\frac{{45}}{360})"


"r= \frac{360}{45 \times2}= 4"


c.

"l=" "2 \times \pi \times 100\times (\frac{{135}}{360}" )


"l=" 75"\pi"